Circular No. 8007

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002hk IN MCG +07-15-6 M. Ganeshalingam and W. Li, University of California at Berkeley

(UCB), report the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered KAIT images taken on Nov. 2.5 (mag ∼ 18.2) and 4.5 UT (mag ∼ 17.9). SN 2002hk is located at α = 7^{h}07^{m}14^{s}.05, δ = +44^{o}48′46′′.2 (equinox 2000.0), which is 11′′.9 east and 14′′.0 south of the nucleus of MCG +07-15-6. A KAIT image taken on Oct. 22.5 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVAE 2002hf AND 2002hh A. V. Filippenko, R. J. Foley, and B. Swift, UCB, report that inspec-

tion of CCD spectra (range 310–1000 nm), obtained on Nov. 2 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2002hf (IAUC 8004) is of type Ic, within one week past maximum brightness. Similarly, spectra of SN 2002hh (IAUC 8005) show it to be a very young, highly reddened type-II supernova; broad, very-low-contrast Hα emission and absorption are visible, but the continuum is nearly featureless and very red. Strong, narrow (interstellar) Na I D absorption is present.

SUPERNOVA 2002hg IN NGC 3306 R. Chornock, S. Jha, and A. V. Filippenko, UCB; and B. Barris, Uni-

versity of Hawaii, report that a spectrum (range 400–1000 nm), obtained on Nov. 1.6 UT with the Keck II 10-m telescope (+ ESI), of SN 2002hg (cf. IAUC 8004) shows it to be a type-II supernova after maximum light, with well-developed P-Cyg lines of the H Balmer series, Fe II, and Ca II. Adopting the NASA/IPAC Extragalactic Database recession velocity for NGC 3306 of 2889 km/s yields an expansion velocity for the supernova of 4900 km/s, measured from the trough of the Hβ profile.

G. Pignata and F. Patat, European Southern Obsrvatory; and M. Turatto, Istituto Nazionale di Astrofisica, Padua, on behalf of the European RTN (cf. IAUC 7961, 7987), report that preliminary reduction of a spectrum (range 350–880 nm; resolution 1.1 nm) of SN 2002hg, obtained at Calar Alto on Nov. 2.22 UT with the 2.2-m reflector (+ CAFOS), shows that this object is a typical type-II supernova a few weeks after maximum light, with strong Balmer lines of H, Fe II, Ca II, and Na I with well-developed P-Cyg profiles. The expansion velocity of the photosphere, deduced from the Hα absorption minimum, is 6800 km/s.

2002 November 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green